OpenStudy (anonymous):

2x−6=8x+3 I'm haveing trouble with this can someone please help me

OpenStudy (anonymous):

yes.it is easy 2x-6=8x+3 2x-8x=3+6 -6x=9 x=-1.5

OpenStudy (anonymous):

can you describe how you did please?

OpenStudy (animelover91804):

Simplifying 3 + -2x = -6 + -8x Solving 3 + -2x = -6 + -8x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '8x' to each side of the equation. 3 + -2x + 8x = -6 + -8x + 8x Combine like terms: -2x + 8x = 6x 3 + 6x = -6 + -8x + 8x Combine like terms: -8x + 8x = 0 3 + 6x = -6 + 0 3 + 6x = -6 Add '-3' to each side of the equation. 3 + -3 + 6x = -6 + -3 Combine like terms: 3 + -3 = 0 0 + 6x = -6 + -3 6x = -6 + -3 Combine like terms: -6 + -3 = -9 6x = -9 Divide each side by '6'. x = -1.5 Simplifying x = -1.5

OpenStudy (anonymous):

yes of course . we need to put x on one side , but number on another we have 2x on left side but 8x on right , it means that we need to change +8x on -8x and it will be on left side 2x-8x but on right side we need have numbers for 3 we add 6 because it had minus but now we change for minus

OpenStudy (anonymous):

excuse me for plus

OpenStudy (anonymous):

thank you so much for your help.

OpenStudy (animelover91804):

Sorry if i put to much, Heh

OpenStudy (anonymous):

but you understand?

OpenStudy (anonymous):

oh your all right

OpenStudy (animelover91804):

I hope you understand it

OpenStudy (anonymous):

I do. But I have another thing I'm stuck on.

OpenStudy (anonymous):

The expression 62.4d-21062.4d−21062, point, 4, d, minus, 210 gives the number of Indian rupees you receive when you exchange \$d$ddollar sign, d at the local currency exchange. How many Indian rupees do you receive when you exchange \$13$13dollar sign, 13?

OpenStudy (anonymous):

thanks for the medal lol

OpenStudy (anonymous):

can you also one time type what you want , because i nothing understand

OpenStudy (anonymous):

about inadian rupees

OpenStudy (anonymous):

huh?